# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ V.2 Higgs=SPT
An SPT phase is a state of matter that cannot be adiabatically connected to a trivial phase under local symmetry-preserving perturbations without closing a gap. In general, a trivial phase can be reached without gap closure only if those symmetries are broken. A classic example of an SPT phase in an interacting lattice model is the spin one-half chain—the cluster model—whose nontrivial topology is protected by $\mathbbm{Z}_{2}\times\mathbbm{Z}_{2}$ symmetry. A consequence of the topological nature of the phase is the presence of gapless states localized at the boundary of an open chain. In [18] it was realized that the cluster SPT order emerges in the Higgs phase of the $\mathbbm{Z}_{2}$-gauged Ising chain in the absence of the electric string tension term. Moreover, the fermionic SPT order of the gauged Kitaev chain in the Higgs regime was identified.
Many 2+1 dimensional SPT phases are known and one of the latest to be added to the inventory of phases lives in the phase diagram of the $2+1$ dimensional $\mathbbm{Z}_{2}$ gauge theory coupled minimally to Ising matter. Using the notation of this paper, where the gauge coupling is denoted by $\beta$ and the matter coupling by $\kappa$, the famous deconfined toric code model lives at $\kappa=0$ and $\beta\rightarrow\infty$. It was argued in Ref. [19] that deep in the Higgs limit $\kappa\rightarrow\infty$ and $\beta\rightarrow\infty$ (where both gauge and matter degrees of freedom are frozen in the bulk and the theory has a magnetic 1-form Ising symmetry) the model maps to the two-dimensional cluster SPT protected by the $1-$form magnetic symmetry and a 00-form Ising matter symmetry 101010 For earlier work on SPTs protected by generalized symmetries see for example Refs. [67, 76] . The Ising matter symmetry is trivialized in the bulk by gauging, but survives as a global symmetry in presence of a boundary. This then has gapless surface states localized on symmetry-preserving boundaries. It was further shown that this phase persists to large but finite $\kappa$ at $\beta\to\infty$, where both protecting symmetries remain intact. Going to finite $\beta$ breaks the magnetic $1$-form symmetry yet at $\cramped{\kappa\to\infty}$ only the boundary degrees of freedom can fluctuate and are governed by an effective one-dimensional transverse field Ising model that is in the spontaneously broken phase for large $\beta$. The gapless edge modes thus originate from the Ising degeneracy of this phase. It is expected [77] that the $1-$form symmetry survives in the infrared even though it is explicitly broken away from $\beta\to\infty$. In other words one can think about it as a low-energy emergent symmetry. Protected by the Ising matter 00-form symmetry and the emergent magnetic $1$-form symmetry, the low-energy edge modes thus should survive within some window of parameters even away from $\beta\to\infty$ limit. It was verified numerically that there is indeed a phase with a gapless boundary within the entire Higgs regime of the model. The phase diagram of this model bears a great deal of resemblance to our Fig. 1 with a phase transition line originating at a triple point at the innermost corner of the deconfined phase and ending on a critical endpoint. Within the resolution of the DMRG numerics of Ref. [19] the phase boundary to the boundary gapless phase extends from the critical endpoint.
A follow-up work extended the results to $U(1)$ Higgs phases [20]. In 3+1D the relevant symmetries are the matter U($1$) 00-form symmetry associated with $q=1$ charge conservation and an exact magnetic $U(1)$ $1$-form symmetry, whose conserved current $J_{\rm mag}\equiv da$. We may couple the currents to background gauge fields—one-form $A$ for the matter current and two-form $B$ for $J_{\rm mag}$. Deep in the Higgs regime, the topological response was found to be
$S_{\rm SPT}=\frac{1}{2\pi}\int B\wedge dA.$ (117)
@@
This SPT response is a consequence of a mixed anomaly between the matter and magnetic symmetries, which is cured by adding boundary degrees of freedom. For the $3+1$D Higgs phase the boundary theory takes the form
[…]
@@
for the conjugate pair: a compact scalar field $\varphi$ and a compact $U(1)$ gauge field $\vartheta$. This describes a boundary $U(1)$ superfluid phase. The continuum discussion was supplemented in Ref. [20] with a lattice Villain formulation, where monopoles are under control and the magnetic $1$-form symmetry is preserved.
We have briefly reviewed the results of [18, 19, 20] on connections between SPT and Higgs phases. But what bearing, if any, do the results we have reported have on this scenario? After all, at no stage of this work did we invoke bulk topological arguments to understand the existence of the boundary phase transition. It was possible to understand the global symmetry breaking purely at the level of the boundary theory. Even so, given the arguments sketched above, one might seek a description in terms of an SPT order since it would shed light on how robust the boundary symmetry breaking actually is.
In the 4D compact Wilson $U(1)$ model coupled to the fundamental Higgs discussed in Sec. II the 1-form $U(1)$ magnetic symmetry is broken explicitly by monopoles. However, in the limit $\beta\to\infty$, where electromagnetic fields freeze, monopole creations/annihilations are suppressed and the magnetic 1-form symmetry emerges. Based on general arguments of [20], the interplay of the 0-form matter and 1-form magnetic $U(1)$ symmetries should give rise to a gapless symmetry-preserving boundary. Sure enough we have established the presence of such a boundary phase numerically. Moreover, in agreement with arguments of [19, 20], the gapless nature of the boundary survives beyond the symmetric $\beta\to\infty$ limit, as is clear in the $\kappa\rightarrow\infty$ limit where the boundary maps to a three dimensional XY model. This much follows from ideas put forward in [19, 20]. However, in contrast to the models studied in [19, 20], our bulk model in the $\beta\rightarrow\infty$ limit becomes gapless for large $\kappa$ since it can be thought of (in the unitary gauge) as the ordered phase of a four-dimensional XY model.111111 In contrast, approaching $\beta\rightarrow\infty$ from the $\kappa\rightarrow\infty$ limit apparently preserves a finite gap. Obviously, it cannot support a gapped SPT phase. It may be interesting to explore this regime of the phase diagram more closely and determine whether it exhibits a gapless SPT order. It is also important and interesting to understand better the role of the $\alpha$ coupling from the point of view of the bulk SPT.
@@
The physics of the non-Abelian gauge theories with Higgs matter gives rise to another puzzle. In Sec. III we have argued that in the limit $\beta,\kappa\gg 1$ boundary is expected to be gapless rather generically. If the Higgs=SPT scenario is general enough to encompass the non-Abelian cases too, one should be able to identify the protecting (higher-form) bulk symmetries. Pure $\mathrm{SU}(N)$ gauge theory has a $\mathbbm{Z}_{N}$ center 1-form symmetry and the electric field (Wilson) lines in $\mathrm{SU}(N)$ transform non-trivally under the center of the group. The center symmetry, however, is broken explicitly in the presence of the fundamental Higgs matter. This should be analogous to the breaking of electric one-form symmetry in the Abelian case discussed above. But, in the limit $\beta\to\infty$ the Abelian theory also has a magnetic $(D-3)$-form symmetry which, together with the matter symmetry, protects the SPT order. In contrast, ’t Hooft lines transform trivially under the center of the group. Currently it is unclear to us whether there is a generalized symmetry present at $\beta\rightarrow\infty$ for non-Abelian gauge group that is an analog of the magnetic one-form symmetry emerging in the Abelian case. Whether SPT order is responsible for the gapless nature of the symmetry-broken boundary phase in the non-Abelian gauge theories coupled to Higgs matter is an open question.
Taking a broader perspective, it would be interesting to search for mutual anomalies between the bulk matter (equivalently, boundary flux) symmetry and any emergent higher-form symmetries which may be present in the $\cramped{\beta\to\infty}$ limit arising from suppressing homotopy defects of the gauge field.
Acknowledgements.
@@
PM and PR acknowledge useful discussions with Pedro Bicudo and Nuno Cardoso as a part of a related collaboration. KTKC acknowledges Chris Hooley for useful discussion. S.M. is supported by Vetenskapsrådet (grant number 2021-03685), Nordita and STINT. This work was in part supported by the Deutsche Forschungsgemeinschaft under the cluster of excellence ct.qmat (EXC-2147, project number 390858490).
@@ Appendix A Discrete Differential Calculus for Abelian Fields
We utilize notation which mimics continuum differential forms on a lattice, borrowed from algebraic topology [78] and used extensively in Section IV. Pedagogical treatments can be found in [29, 48]. Fields are described (locally) as differential forms (i.e. anti-symmetric tensors), whose primary property is that they can be integrated over surfaces. For our purposes we can think of them simply as functions of such surfaces, i.e. if $\omega$ is a differential $k$-form and $U$ an oriented $k$-dimensional surface, $\omega$ acts on $U$ by integration
$\omega(U):=\int_{U}\omega,$ (119)
@@
$\omega(U_{1}+U_{2})$ $=\omega(U_{1})+\omega(U_{2}),$ (120)
where $-U$ denotes the reversed orientation. Lastly, we can naturally define the derivative of a $k$-form $\omega$ to be a $(k+1)$-form $\mathrm{d}\omega$ whose value on a $(k+1)$-dimensional surface $U$ is defined by Stoke’s theorem,
$\int_{U}\mathrm{d}\omega:=\int_{\partial U}\omega,$ (121)
@@
In this paper, our space(time) is a $D$-dimensional cubical cell complex—a collection of $k$-cells for $k=0,\ldots,D$, i.e. vertices, links, plaquettes, cubes, hypercubes, etc., where $k$-cells are glued along their $(k-1)$-dimensional boundary cells. We denote the collection of all cells by $X$ and the collection of all $k$-cells by $X_{k}$. Because we also consider open boundaries in the form of Fig. 2, we let $X_{k}$ denote just the bulk $k$-cells and $\partial X_{k}$ denote the $k$-cells in the boundary layer which touch the vacuum.
@@
$\partial:\mathcal{C}_{k}\to\mathcal{C}_{k-1},$ (122)
which distributes over the linear combinations on $k$-cells, and sends each oriented $k$-cell to the linear combination of its oriented boundary $(k-1)$-cells.
The sensible lattice analog of a differential form, i.e. a discrete $k$-form, also called a $k$-cochain, is a function which (i) maps chains to numbers, Eq. 119, and (ii) disributes over linear combinations, Eq. 120. In full generality, a discrete $k$-form $\omega$ is a linear map from chains to elements of any Abelian group $\mathcal{G}$,
$\omega:\mathcal{C}_{k}\to\mathcal{G}.$ (123)
In practice what this means is that a disrete $k$-form is defined by its value on each oriented $k$-cell $c$, and satsfies $\cramped{\omega(-c)=-\omega(c)}$, where $-\omega$ is understood as the inverse operation in the Abelian group $\mathcal{G}$. The space of discrete $k$-forms is denoted $\mathcal{C}^{k}(\mathcal{G})$. Since we already have a natural notion of the boundary operation on chains, we can define a natural exterior derivative operation on cochains, $\cramped{\mathrm{d}:\mathcal{C}^{k}\to\mathcal{C}^{k+1}}$, according to Stoke’s theorem, i.e.
$\mathrm{d}\omega(U):=\omega(\partial U),$ (124)
@@
$(u,w)=\sum_{c}u_{c}w_{c}.$ (126)
Using this, we can define an adjoint of the boundary operator, which we call the coboundary,121212 Note that this differs from the algebraic topology terminology, where the discrete exterior derivative is often called the coboundary.
$(u,\partial v)=(\partial^{\dagger}u,v).$ (127)
@@
$\partial^{\dagger}c=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(\partial^{\dagger}c)_{c^{\prime}}c^{\prime}=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(\partial^{\dagger}c,c^{\prime})c^{\prime}=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(c,\partial c^{\prime})c^{\prime}$ (128)
By choosing to orient all the $c^{\prime}$ in the sum such that $(c,\partial c^{\prime})=+1$ or 00, we can read this to say that the coboundary of an oriented $k$-cell $c$ is the sum of all oriented $(k+1)$-cells containing $+c$ in their positively oriented boundary. For example, the coboundary of a point $i$ is the set of oriented links which terminate at it, the coboundary of a link $\ell$ is the set of plaquettes $p$ touching it, oriented so that $\partial p$ circulates in the same direction as $\ell$ is oriented, etc.
@@
$\mathrm{d}^{\dagger}\omega(c):=\omega(\partial^{\dagger}c),$ (129)
@@
$\langle\alpha,\beta\rangle=\sum_{c\in X_{k}}\alpha(c)^{*}\beta(c)$ (130)
@@
$=\sum_{\mathclap{c^{\prime}\in X_{k-1}}}\alpha(\partial^{\dagger}c^{\prime})^{*}\,\beta(c^{\prime})=\langle\mathrm{d}^{\dagger}\alpha,\beta\rangle.$ (131)
Figure 12: (a,b) Phase diagram of the 4D $U(1)$ Abelian Higgs model for $L=16$ revealed through (a) expectation of the minimal Wilson loop, $\langle 1-\mathrm{Re}\,W_{p}\rangle$, and (b) expectation of the Wilson link, $\mathrm{Re}\,\Lambda_{\ell}\rangle$, averaged over all plaquettes and links, respectively. The color ranges from $0.1$ (black) to $0.8$ (yellow). (c) Cuts of $\langle 1-\mathrm{Re}\,W_{p}\rangle$, at constant $\beta$ for different values of $\beta$ from $0.525$ to $1.000$ in steps of $0.025$ from right to left. The system is isotropic with $L=16$ and $50000$ sweeps. The data shows a clear sign of a phase transition for larger $\beta$ and a crossover for smaller $\beta$. The data is consistent with the presence of a first order transition with a critical endpoint at about $\beta=0.85$.
Lastly, we review a bit of basic algebraic topology terminology used in Section IV. The boundary operation is nilpotent, $\partial^{2}=0$, and thus defines an exact sequence of maps,
$0\to\mathcal{C}_{D}\xrightarrow{\partial_{D}}\cdots\mathcal{C}_{1}\xrightarrow{\partial_{1}}\mathcal{C}_{0}\to 0$ (132)
This allows us to define the homology groups, $H_{k}:=\mathrm{ker}\,\partial_{k}/\mathrm{im}\,\partial_{k+1}.$ The interpretation of these quotient groups is that they classify the non-contractible $k$-dimensional surfaces. Here, $\mathrm{ker}\,\partial$ is generated by the set of surfaces without boundaries (called cycles), while $\mathrm{im}\,\partial$ is generated by the set of surfaces which are boundaries of a $(k+1)$-dimensional volume, and are therefore contractible. The elements of $H_{k}$ are equivalence classes of surfaces which differ by a boundary, i.e. homology classes.

### v2

@@ 5.2 Higgs = SPT
It would be remiss of us to conclude this paper without a description of some of the work that motivated this study—namely the papers [21, 22, 23] discussing the relationship of Higgs phases to symmetry-protected topological (SPT) phases. We briefly summarize the main findings therein and comment on how our results bear on the generality of this relationship. For more recent works related to this topic see Refs. [108, 109, 110]. An SPT phase is a state of matter that cannot be adiabatically connected to a trivial phase under local symmetry-preserving perturbations without closing a gap. In general, a trivial phase can be reached without gap closure only if those symmetries are broken. A classic example of an SPT phase in an interacting lattice model is the spin one-half chain—the cluster model—whose nontrivial topology is protected by $\mathbbm{Z}_{2}\times\mathbbm{Z}_{2}$ symmetry [111, 112]. A consequence of the topological nature of the phase is the presence of gapless states localized at the boundary of an open chain.
5.2.1 Review of Higgs=SPT
The “Higgs=SPT” connection was first raised in Ref. [21], which considered gauging the $\mathbbm{Z}_{2}$ fermion parity symmetry of the Kitaev chain, which hosts an SPT phase. Gauging trivializes the symmetry in the bulk, but by the same arguments we have discussed throughout this paper, it is instead realized non-trivially on the boundary. In the Higgs regime a fermionic SPT order emerges (that belongs to the same class as a stack of two Kitaev chains) protected by the $\mathbbm{Z}_{2}$ boundary fermion parity symmetry and the $\mathbbm{Z}_{2}$ 0-form magnetic symmetry of the gauge field. Allowing instantons (magnetic tunneling events) explicitly breaks the 0-form magnetic symmetry and kills the SPT. In the similar spirit, it was argued that by gauging the $\mathbbm{Z}_{2}$ parity symmetry of the transverse field Ising model, in the Higgs phase one ends up with the cluster SPT order.
These 1D results were extended in Ref. [22] to the Higgs phases of $\mathbbm{Z}_{2}$ gauge theory with matter in general dimensions with electric boundary conditions (Fig.˜2). The phase diagram is qualitatively the same as Fig.˜1, with (within our notation) the deconfined phase at large $\beta$ and small $\kappa$ exhibiting $\mathbbm{Z}_{2}$ toric code topological order (i.e. 1-form symmetry breaking) [28, 22, 97]. The $\mathbbm{Z}_{2}$ magnetic symmetry is a $\cramped{(d-1)}$-form symmetry.121212 Note that the magnetic symmetry is (d1)(d-1)-form for discrete gauge group and $\cramped{(d-2)}$-form for $\mathrm{U}(1)$, where $d$ is the dimension of space. The global $\mathbbm{Z}_{2}$ 0-form Ising matter symmetry is trivial in the bulk but is realized at the boundary. It was shown that on the line $\cramped{\beta=\infty}$ and in the large-$\kappa$ Higgs limit there is a cluster SPT order, protected by the $\mathbbm{Z}_{2}$ higher-form magnetic symmetry and the matter symmetry, hosting gapless surface states and a boundary ground state degeneracy.131313 For earlier work on SPTs protected by generalized symmetries see for example Refs. [74, 113] Going to finite $\beta$ explicitly breaks the magnetic $1$-form symmetry, but as long as the magnetic symmetry is higher-form it is expected to survive in the infrared as a low-energy emergent symmetry, because the magnetic defects are gapped at large $\beta$.141414 The magnetic symmetry is equivalent to the closure of magnetic flux surfaces. The presence of magnetic defects allows the fluxes to end without closing, and thus explicitly breaks the symmetry. At large $\beta$ the magnetic defects will have a large gap, and therefore at low energies magnetic fluxes will again be closed. Thus the symmetry is expected to be emergent in the infrared [114]. Indeed this argument explains why the deconfined phase, which spontaneously breaks the higher-form electric and magnetic symmetries, is stable in the presence of gapped magnetic and electric charges [4]. Moreover, it was shown that in the $\cramped{\kappa=\infty}$ limit the boundary degrees of freedom are governed by an Ising model charged under the global matter symmetry, which is in the spontaneously broken phase for large $\beta$. The boundary ground state degeneracy of the putative $\cramped{\beta=\infty}$ SPT were directly identified with the Ising degeneracy of this phase. The boundary symmetry breaking was verified numerically with a phase diagram qualitatively the same as Fig.˜1, and within the resolution of the DMRG numerics of Ref. [22] the phase boundary to the boundary gapless phase extends from the critical endpoint. However, one cannot simply identify the boundary-symmetry-broken regime with a bulk SPT, indeed Ref. [22] observed that the position of the transition depends on microscopic details near the boundary (such as the deformation $\alpha$ which we have considered). Instead it was argued that certain properties of SPT (such as edge modes and degeneracies of the entanglement spectrum) should be stable in the Higgs regime in an open region of the phase diagram in the vicinity of the $\cramped{\beta\to\infty}$ limit.151515 The entanglement spectrum in this model was investigated recently in [115]. In particular, the mutual anomaly between the matter and magnetic symmetries at the boundary is robust, and correspondingly the edge mode degeneracies should be as well until a gap is closed.
These ideas were extended in Ref. [23], which considers $\mathrm{U}(1)$ Higgs models with exact magnetic symmetry, i.e. without magnetic monopoles. Starting from the continuum theory, Eq.˜1, the mutual anomaly can be exposed by coupling to background gauge fields for both symmetries—1-form $A_{\text{mat}}$ for the matter symmetry and $\cramped{(d-1)}$-form $B_{\text{mag}}$ for the magnetic symmetry. Deep in the Higgs regime, the bulk topological response was found to be
$S_{\rm SPT}=\frac{1}{2\pi}\int B_{\text{mag}}\wedge\mathrm{d}A_{\text{mat}}.$ (119)
@@ 5.2.1 Review of Higgs=SPT
For a system with open boundaries this SPT response is not gauge invariant without adding appropriate boundary degrees of freedom which cancel the anomaly. For the 3+13+1D Higgs phase, the boundary theory takes the form
$S_{\rm Boundary}=\frac{1}{2\pi}\int\mathrm{d}\varphi\wedge\mathrm{d}\vartheta$ (120)
@@
for the conjugate pair: a compact scalar field $\varphi$ and a compact $U(1)$ gauge field $\vartheta$. This describes a boundary $U(1)$ superfluid phase, which has the correct mixed anomaly to cure the bulk one [6]. This continuum discussion was supplemented with a lattice Villain formulation, where monopoles are under control and the magnetic $1$-form symmetry is preserved.
As it stands, we can summarize the Higgs=SPT situation as follows. For Abelian gauge fields coupled to fundamental matter with exact magnetic symmetry, the matter symmetry is realized at the boundary and shares a mutual anomaly with the bulk magnetic symmetry. The bulk of the system can be identified as an SPT protected by these two symmetries, and the SPT boundary modes originate from the spontaneously broken matter symmetry: for $\mathbbm{Z}_{N}$ gauge group these are robust boundary discrete degeneracies, and for U(1) gauge group these are gapless edge modes identified with the boundary Goldstone modes. The authors of both Ref. [22] and Ref. [23] claim that some aspects of the SPT, in particular the edge modes and entanglement spectrum degeneracies, should be robust to explicit breaking of magnetic symmetry (when it is a higher-form symmetry), even if the bulk is no longer strictly an SPT (and can indeed be continuously connected to the trivial confined phase).
5.2.2 Insights, Challenges, and Future Directions
Our work potentially extends the Higgs=SPT story, but also provides new challenges for it. Firstly, the general Higgs=SPT arguments generalize directly to the higher-form Abelian models we considered in Section˜4 if magnetic symmetry is enforced exactly. Secondly, we can say that spontaneous breaking of global charge conservation symmetry at the boundary is a universal feature of the Higgs regime, and that this statement applies to both Abelian and non-Abelian models. This could indicate that the Higgs-SPT relation holds in a much broader set of cases than originally envisioned. However, it may also be that the boundary symmetry breaking is a broader phenomenon than the Higgs=SPT relation, which happens to coincide with the SPT edge modes when the relation holds.
We wish to point out an open issue for the Higgs=SPT connection with continuous gauge groups in the models we have considered here. In the Abelian-Higgs model, when the gauge group is discrete the gapped Higgs phase is continuously connected to the $\cramped{\beta=\infty}$ limit where the magnetic symmetry is exact and the ground state is an SPT [22]. However, there is a subtlety about the large-$\beta$ limit in the U(1) case: the $\cramped{\beta=\infty}$ limit of the Higgs regime is gapless, namely it is gauge-equivalent to a bulk superfluid. Therefore, if there is an SPT in the U(1) Fradkin-Shenker model at $\cramped{\beta=\infty}$ where the magnetic symmetry is exact it must be gapless, and is therefore not adiabatically connected to the gapped Higgs phase when the gauge coupling is turned on.161616 This same problem with the $\beta=\infty$ limit is also present in the Villainized model studied by Ref. [23] (in contrast to the Wilson model we have studied), but in that model one can effectively tune the magnetic monopole mass to infinity while holding $\beta$ finite to obtain an exact magnetic symmetry. Furthermore, the limiting point $\cramped{\beta=\kappa=\infty}$ in the phase diagram is problematic, because the gap scales as $\kappa/\beta$ as one approaches it [14]: approaching from $\cramped{\beta=\infty}$ the bulk is gapless, while approaching from $\cramped{\kappa=\infty}$ the bulk is gapped.171717 This limit was key to establishing the fixed-point SPT Hamiltonian in the $\mathbbm{Z}_{2}$ case in Ref. [22]. It is therefore unclear that the boundary edge modes of the gapped Higgs regime discussed in Section˜2 and Section˜4, particularly in the $\cramped{\kappa=\infty}$ limit, can be identified with those of a putative $\cramped{\beta=\infty}$ SPT. Clearly this same issue holds for any continuous gauge group, in particular to the non-Abelian cases studied in Section˜3.
@@ 5.2.2 Insights, Challenges, and Future Directions
The physics of the non-Abelian gauge groups gives rise to a further puzzle. In Section˜3 we have argued that in the limit $\beta,\kappa\gg 1$ the boundary is generically expected to be gapless since it breaks a continuous symmetry. If the Higgs=SPT scenario is general enough to encompass the non-Abelian cases too, one should be able to identify the protecting (higher-form) bulk “magnetic” symmetry. Pure $\mathrm{SU}(N)$ gauge theory has no magnetic symmetry since the fundamental group is trivial, and it is unclear whether there is a generalized symmetry present at $\cramped{\beta=\infty}$ for non-Abelian gauge groups that is an analog of the Abelian magnetic symmetry.
Let us take a broader perspective and focus on the mutual anomaly which is key to the Higgs=SPT phenomenology. The core idea in the Abelian theories is as follows. Given an Abelian $k$-form symmetry $\mathcal{G}$ there is a corresponding “dual” $\cramped{(d-k-n)}$-form $\mathcal{G}$ symmetry corresponding to suppressing defects in the ordered phase, where $n=0$ (11) if $\mathcal{G}$ is discrete (continuous), and these share a mutual anomaly [22, 23, 114].181818 For example: a 0-form $\mathbbm{Z}_{n}$ symmetry has a corresponding $d$-form $\mathbbm{Z}_{n}$ symmetry associated to suppression of domains walls (conservation of symmetry breaking sector); a 0-form U(1) symmetry has a corresponding (d1)(d-1)-form U(1) symmetry associated to suppression of vortices (conservation of winding number); and a 1-form U(1) symmetry has a corresponding (d2)(d-2)-form U(1) symmetry associated to suppression of monopoles (i.e. magnetic symmetry, conservation of magnetic flux). When the symmetry is gauged, it is realized as a $k$-form symmetry on the $\cramped{(d-1)}$-dimensional boundary, while the bulk has a new $\cramped{((d-1)-k-n)}$-form magnetic symmetry, which we have emphasized has exactly the right degree to play the role of the dual symmetry on the boundary, and indeed shares the correct mutual anomaly [22, 23]. Some generalization of this mechanism should likely also hold true for gauging 0-form non-Abelian symmetries and spontaneously breaking them at the boundary. In general the emergent symmetries arising from spontaneously breaking a 0-form $\mathcal{G}$ symmetry do not form a group but rather a symmetry category, and these symmetries share a mixed anomaly with $\mathcal{G}$ [114, 116]. Upon gauging the 0-form $\mathcal{G}$ symmetry, forcing it to the boundary, one expects that the resulting gauge field has a “magnetic” symmetry category which preserves the correct anomaly structure at the boundary. The precise mathematical structure of the resulting mutual anomaly after gauging is likely complicated due to the fact that the boundary symmetry is generated by the total charge (gauge plus matter, cf. Section˜3.1.3), rather than just the charge of the Higgs field. However, we believe that identifying the relevant generalized symmetries of the non-Abelian gauge field and matching the boundary anomaly are the required ingredients to extend the Higgs-SPT connection to the non-Abelian setting.
Acknowledgements
@@
PM and PR acknowledge useful discussions with Pedro Bicudo and Nuno Cardoso as a part of a related collaboration. KTKC acknowledges Chris Hooley for useful discussion.
Funding information
S.M. is supported by Vetenskapsrådet (grant number 2021-03685), Nordita and STINT. This work was in part supported by the Deutsche Forschungsgemeinschaft under the cluster of excellence ct.qmat (EXC-2147, project number 390858490).
Data Availability
Data and code used in this work are available upon request.
@@ Appendix A Discrete Differential Calculus for Abelian Fields
We utilize notation which mimics continuum differential forms on a lattice, borrowed from algebraic topology [117] and used extensively in Section˜4. Pedagogical treatments can be found in [34, 56]. Fields are described (locally) as differential forms (i.e. anti-symmetric tensors), whose primary property is that they can be integrated over surfaces. For our purposes we can think of them simply as functions of such surfaces, i.e. if $\omega$ is a differential $k$-form and $U$ an oriented $k$-dimensional surface, $\omega$ acts on $U$ by integration
$\omega(U):=\int_{U}\omega,$ (A.1)
@@
$\omega(U_{1}+U_{2})$ $=\omega(U_{1})+\omega(U_{2}),$ (A.2)
where U-U denotes the reversed orientation. Lastly, we can naturally define the derivative of a $k$-form $\omega$ to be a $(k+1)$-form $\mathrm{d}\omega$ whose value on a $(k+1)$-dimensional surface $U$ is defined by Stoke’s theorem,
$\int_{U}\mathrm{d}\omega:=\int_{\partial U}\omega,$ (A.3)
@@
In this paper, our space(time) is a $D$-dimensional cubical cell complex—a collection of $k$-cells for $k=0,\ldots,D$, i.e. vertices, links, plaquettes, cubes, hypercubes, etc., where $k$-cells are glued along their (k1)(k-1)-dimensional boundary cells. We denote the collection of all cells by $X$ and the collection of all $k$-cells by $X_{k}$. Because we also consider open boundaries in the form of Fig.˜2, we let $X_{k}$ denote just the bulk $k$-cells and $\partial X_{k}$ denote the $k$-cells in the boundary layer which touch the vacuum.
@@
$\partial:\mathcal{C}_{k}\to\mathcal{C}_{k-1},$ (A.4)
which distributes over the linear combinations on $k$-cells, and sends each oriented $k$-cell to the linear combination of its oriented boundary (k1)(k-1)-cells.
The sensible lattice analog of a differential form, i.e. a discrete $k$-form, also called a $k$-cochain, is a function which (i) maps chains to numbers, Eq.˜A.1, and (ii) disributes over linear combinations, Eq.˜A.2. In full generality, a discrete $k$-form $\omega$ is a linear map from chains to elements of any Abelian group $\mathcal{G}$,
$\omega:\mathcal{C}_{k}\to\mathcal{G}.$ (A.5)
In practice what this means is that a disrete $k$-form is defined by its value on each oriented $k$-cell $c$, and satsfies $\cramped{\omega(-c)=-\omega(c)}$, where $\omega$ is understood as the inverse operation in the Abelian group $\mathcal{G}$. The space of discrete $k$-forms is denoted $\mathcal{C}^{k}(\mathcal{G})$. Since we already have a natural notion of the boundary operation on chains, we can define a natural exterior derivative operation on cochains, $\cramped{\mathrm{d}:\mathcal{C}^{k}\to\mathcal{C}^{k+1}}$, according to Stoke’s theorem, i.e.
$\mathrm{d}\omega(U):=\omega(\partial U),$ (A.6)
@@
$(u,w)=\sum_{c}u_{c}w_{c}.$ (A.8)
Using this, we can define an adjoint of the boundary operator, which we call the coboundary,191919 Note that this differs from the algebraic topology terminology, where the discrete exterior derivative is often called the coboundary.
$(u,\partial v)=(\partial^{\dagger}u,v).$ (A.9)
@@
$\partial^{\dagger}c=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(\partial^{\dagger}c)_{c^{\prime}}c^{\prime}=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(\partial^{\dagger}c,c^{\prime})c^{\prime}=\sum_{\mathclap{c^{\prime}\in X_{k+1}}}(c,\partial c^{\prime})c^{\prime}$ (A.10)
By choosing to orient all the $c^{\prime}$ in the sum such that $(c,\partial c^{\prime})=+1$ or 0, we can read this to say that the coboundary of an oriented $k$-cell $c$ is the sum of all oriented $(k+1)$-cells containing +c+c in their positively oriented boundary. For example, the coboundary of a point $i$ is the set of oriented links which terminate at it, the coboundary of a link $\ell$ is the set of plaquettes $p$ touching it, oriented so that $\partial p$ circulates in the same direction as $\ell$ is oriented, etc.
@@
$\mathrm{d}^{\dagger}\omega(c):=\omega(\partial^{\dagger}c),$ (A.11)
@@
$\langle\alpha,\beta\rangle=\sum_{c\in X_{k}}\alpha(c)^{*}\beta(c)$ (A.12)
@@
$=\sum_{\mathclap{c^{\prime}\in X_{k-1}}}\alpha(\partial^{\dagger}c^{\prime})^{*}\,\beta(c^{\prime})=\langle\mathrm{d}^{\dagger}\alpha,\beta\rangle.$ (A.13)
Qualitatively, one may think of the differential as a generalized gradient, describing how a $k$-form varies across $(k+1)$-cells, and the codifferential as a generalized divergence, describing how a $k$-form “flows into” (k1)(k-1)-cells.
[…]
Figure 13: (a,b) Phase diagram of the 4D $U(1)$ Abelian Higgs model for $L=16$ revealed through (a) expectation of the minimal Wilson loop, $\langle 1-\mathrm{Re}\,W_{p}\rangle$, and (b) expectation of the Wilson link, $\langle\mathrm{Re}\,\Lambda_{\ell}\rangle$, averaged over all plaquettes and links, respectively. The color ranges from $0.1$ (black) to $0.8$ (yellow). (c) Cuts of $\langle 1-\mathrm{Re}\,W_{p}\rangle$, at constant $\beta$ for different values of $\beta$ from $0.525$ to $1.000$ in steps of $0.025$ from right to left. The system is isotropic with $L=16$ and $50000$ sweeps. The data shows a clear sign of a phase transition for larger $\beta$ and a crossover for smaller $\beta$. The data is consistent with the presence of a first order transition with a critical endpoint at about $\beta=0.85$.
Lastly, we review a bit of basic algebraic topology terminology used in Section˜4. The boundary operation is nilpotent, $\partial^{2}=0$, and thus defines an exact sequence of maps,
$0\to\mathcal{C}_{D}\xrightarrow{\partial_{D}}\cdots\mathcal{C}_{1}\xrightarrow{\partial_{1}}\mathcal{C}_{0}\to 0.$ (A.14)
We can define the homology groups, $\cramped{H_{k}:=\mathrm{ker}\,\partial_{k}/\mathrm{im}\,\partial_{k+1}.}$ The interpretation of these quotient groups is that they classify the non-contractible $k$-dimensional surfaces. Here, $\mathrm{ker}\,\partial$ is generated by the set of surfaces without boundaries (called cycles), while $\mathrm{im}\,\partial$ is generated by the set of surfaces which are boundaries of a $\cramped{(k+1)}$-dimensional volume, and are therefore contractible. The elements of $H_{k}$ are equivalence classes of surfaces which differ by a boundary, i.e. homology classes.